# Scaling Gaussian Process Regression with Full Derivative Observations

### v1

@@ Appendix C Background on DDSVGP and DSKI
Variational inducing points. An inducing point method (Snelson & Ghahramani, 2005; Quinonero-Candela & Rasmussen, 2005) introduces a set of $m\ll n$ inducing points $\mathbf{z}=(z_{i}\in\mathbb{R}^{d})_{i=1}^{m}$ and associated inducing variables $\mathbf{u}=(u_{i}\in\mathbb{R})_{i=1}^{m}$ as a proxy for the given inputs $\mathbf{x}$ and outputs $\mathbf{y}$ respectively. In an inducing point method such as Sparse Variational Gaussian Processes (SGPR) (Titsias, 2009) and Stochastic Variational Gaussian Process (SVGPs) (Hensman et al., 2013), the inducing points and variables are related to the dataset as
@@
A SVGP can be extended to the setting with derivatives, i.e., a DSVGP (Padidar et al., 2021). More concretely, define an inducing variable $\tilde{u}_{i}=\begin{pmatrix}u_{i}&\nabla u_{i}^{T}\end{pmatrix}^{T}$ that models the value and its gradient at an inducing point $z_{i}$. A DSVGP introduces learnable parameters $\mathbf{m}^{(m\times 1)}$, $d\mathbf{m}^{(dm\times 1)}$, $\mathbf{S}^{(m\times m)}$, $d\mathbf{S}^{(1\times m)}$, and $d^{2}\mathbf{S}^{(dm\times dm)}$ to define a variational posterior $q(\tilde{\mathbf{u}})=\mathcal{N}(\tilde{\mathbf{u}}\,|\,\tilde{\mathbf{m}},\tilde{\mathbf{S}})$ where
@@
Structured kernel interpolation. One problem with inducing point methods is the requirement that $m\ll n$. SKI (Wilson & Nickisch, 2015) overcomes this limitation by using the approximate kernel
@@
The posterior is tractable to compute since the lattice structure given by the interpolation points enables fast matrix-vector multiplies (MVMs) so that conjugate gradient (CG) methods can be used to solve large systems of linear equations. The complexity of a single MVM is $O(n4^{d}+m\log m)$ with a cubic interpolation function (Wilson & Nickisch, 2015) where $m$ is the number of interpolation points. However, the asymptotic dependence on $d$ limits the application of SKI to low dimensions.
DSKI (Eriksson et al., 2018) extends the SKI (Wilson & Nickisch, 2015) method to create a scalable GP regression method that handles derivative information. To accomplish this, DSKI uses the following approximate kernel

### v2

@@ Appendix E Additional Background on DDSVGP and DSKI
Variational inducing points. An inducing point method (Snelson and Ghahramani, 2005; Quinonero-Candela and Rasmussen, 2005) introduces a set of $m\ll n$ inducing points $\mathbf{z}=(z_{i}\in\mathbb{R}^{d})_{i=1}^{m}$ and associated inducing variables $f(\mathbf{z})=\mathbf{u}=(u_{i}\in\mathbb{R})_{i=1}^{m}$ as a proxy for the given inputs $\mathbf{x}$ and outputs $\mathbf{y}$ respectively. In an inducing point method such as Sparse Variational Gaussian Processes (SGPR) (Titsias, 2009) and Stochastic Variational Gaussian Process (SVGPs) (Hensman et al., 2013), the inducing points and variables are related to the dataset as
@@
[…]
@@
Structured kernel interpolation. One problem with inducing point methods is the requirement that $m\ll n$. SKI (Wilson and Nickisch, 2015) overcomes this limitation by using the approximate kernel
@@
The posterior is tractable to compute since the lattice structure given by the interpolation points enables fast matrix-vector multiplies (MVMs) so that conjugate gradient (CG) methods can be used to solve large systems of linear equations. The complexity of a single MVM is $O(n4^{d}+m\log m)$ with a cubic interpolation function (Wilson and Nickisch, 2015) where $m$ is the number of interpolation points. However, the asymptotic dependence on $d$ limits the application of SKI to low dimensions.
DSKI (Eriksson et al., 2018) extends the SKI (Wilson and Nickisch, 2015) method to create a scalable GP regression method that handles derivative information. To accomplish this, DSKI uses the following approximate kernel